OpenStudy (anonymous):

What is 1747 in scientific notation?

OpenStudy (anonymous):

1.747x10^3 You seriously just need to learn to the rules. Its a super easy concept.

OpenStudy (theeric):

Scientific notation just expresses a number like that. You are really not changing the number. Look! I can show it by multiplying both sides by \(1=\dfrac{10^3}{10^3}\)! \(1747\times\dfrac{10^3}{10^3}=\dfrac{1747}{10^3}\times 10^3=1.747\times 10^3\) And the number is the same, since I just multiplied by \(1\). To find what to multiply, moving the decimal to the left is a positive exponent, moving it to the right is a negative exponent. It's a lot like counting. \(1747\times 10^0\\=174.7\times10^1\\=17.47\times10^2\\=1.747\times 10^3\) See the decimal moving to the left, and the coefficient gaining \(1\) each time? Good luck with any others you have to do!